OpenStudy (abbycross167):

20. When we have zero on the bottom of the slope, we call that no slope because of the rule that zero can never be in the denominator of a fraction. An example of this would be a line that goes through the points (2, -3) and (2, 4). If you plotted these points, what would this graph look like? What should this equation be? My Answer~ M=4+3/2-2 M=7/0 The slope is undefined The line would look like a vertical line. What my teacher said~ one more question here that needs to be answered. What is that question? @atrey6s3x6

OpenStudy (abbycross167):

@atreyu6s3x6

OpenStudy (mathstudent55):

You are correct: 1) The sloe is undefined because it involves division by zero. 2) It is a vertical line. 3) You are missing the equation of the line.

OpenStudy (mathstudent55):

You just need to supply 3) above.

OpenStudy (abbycross167):

Ok would you mind helping me find the equation?

OpenStudy (anonymous):

you got this :)

OpenStudy (mathstudent55):

\(\text {20. When we have zero on the bottom of the slope, we call that no slope}\) \(\text{ because of the rule that zero can never be in the denominator of a fraction.}\) \(\text { An example of this would be a line that goes through the points (2, -3) and}\) \(\text{ (2, 4). If you plotted these points, what would this graph look like?} \) \(\color{red}{ What ~should ~this ~equation ~be?} \)

OpenStudy (mathstudent55):

Notice in that the two given points have x = 2, correct? No matter what the y-coordinate of any point is, x is always 2. In other words, there is no restriction on y. The only restriction is that every point has to satisfy the equation x = 2. That is exactly the equation, x = 2.

OpenStudy (mathstudent55):

Similarly, for a horizontal line, the eqation is y = number

OpenStudy (abbycross167):

Give me a second, I'm trying to do the steps

OpenStudy (abbycross167):

ok @mathstudent55 if I show you the steps I did can you check it?

OpenStudy (abbycross167):

@atreyu6s3x6 can you check my steps?

OpenStudy (anonymous):

sure post them :)

OpenStudy (mathstudent55):

You already showed your slope calculation and showed it ends in division by zero. Then you wrote that the slope is undefined. A line with undefined slope is vertical. All you need to add to your work is The equation of the line is x = 2.

OpenStudy (abbycross167):

Points ~~ (2, -3) and (2, 4). (X1, y1) (x2,y2) (x2,y-3) (x2,y4) M= 4-2/-3-2 M= 2/-5 Y=mx+b this is how far i got...

OpenStudy (anonymous):

he gave you the answer :) @abbycross167

OpenStudy (abbycross167):

oh thank you kindly! @mathstudent55